OpenStudy (anonymous):

Consider the line that passes through the point (3, -6) and has a slope of 4. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (anonymous):

When you show the work can you put it all in one reply. It's easier for me. please.

OpenStudy (anonymous):

ok here we go Part 1 once again you are given a point and a slope right? so to find the equation of the line you need to use the "point-slope" formula, because that is the information you have the point - slope formula is \(y-y_1=m(x-x_1)\) you directly substitute the numbers you are given for \(x_1,y_1,m\) in this case they are \(x_1=3,y_1=-6,m=4\) so you write \[y-(-6)=4(x-3)\] or \[y+6=4(x-3)\] end of Part 1

OpenStudy (anonymous):

Part 2 rewrite \(y+6=4(x-3)\) in the slope - intercept form. the slope intercept form in general is \(y=mx+b\) so we need to solve for \(y\) steps are 1) multiply out on the right using the distributive law \[y+6=4(x-3)\] \[y+6=4x-12\] 2) subtract 6 from both sides \[y=4x-12\] end of Part 2

OpenStudy (anonymous):

part 3 rewrite in standard form. standard form is \(ax+by=c\) to this end we use algebra to put the \(x\) and \(y\) terms on one side of the equal sign \[y=4x-18\] \[-4x+y=-18\] or if you prefer, multiply both sides by \(-1\) and write \[4x-y=18\] end of part 3

OpenStudy (anonymous):

Thank you SO SO SO much.

OpenStudy (anonymous):

that is correct, no step is missing. in your first post you were given two incorrect answers. one was that the "slope - intercept" form was \(y-y_1=m(x-x_1)\) which is incorrect. the other was that the answer was \(y=4x-6\) which was also incorrect

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

SO part three would look like this y=4x−18 multiply both sides by -1 4x−y=18

OpenStudy (anonymous):

not quite. to write in "standard form" you need to change \[y=mx+b\] to \[ax+by=c\] in other words put the variables on one side of the equal sign and the constant on the other

OpenStudy (anonymous):

So it would just be 4x−y=18

OpenStudy (anonymous):

part 3 \[y=4x-18\] subtract \(4x\) from both sides get \[-4x+y=-18\] multiply both sides by \(-1\) get \[4x-y=18\] those are the steps

OpenStudy (anonymous):

yup, that is it \[4x-y=18\]

OpenStudy (anonymous):

you got this now, more or less? we can do another if you like

OpenStudy (anonymous):

I think I got it. Thanks again (: Can you help me with a graph?

OpenStudy (anonymous):

I posted it.